A slice of copper is 1.0mm thick and carries a current of 10A. A magnetic field of flux density 1.0T, correctly applied, produces a maximum Hall voltage of 0.6x10^-6V between the edges of the slice. Calculate the number of free charge carriers per unit volume.


1
Expert's answer
2021-04-01T18:34:36-0400

The Hall voltage on the opposite faces:


"V_H=\frac{IB}{nte},\\\space\\
n=\frac{IB}{V_Hte},\\\space\\
n=\frac{10·1}{(0.6·10^{-6})(0.001)(1.602·10^{-19})}=1.04·10^{29}\text{ m}^{-3}"